40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1681     Entry time: Sat Jul 23 01:16:06 2016
Author: awade and Antonio 
Type: DailyProgress 
Category: Mode matching 
Subject: South path: MM solution into south cavity and FI isolation numbers 

A mode matching solution was found for the beam to be coupled into the south cavity. 

With lengths referenced to the first hole in the second table (2815 mm from the laser head)

f = 143.23 mm (PLCX-25.4-64.4-UV-1064) @ 0.148 m

f = 103.20 mm (PLCX-25.4-46.4-UV-1064) @ 0.589 m

f = 171.92 mm (PLCX-25.4-77.3-UV-1066) @ 0.974 m

A small alteration was made as we settled on the final solution, the first lens was shifted back to the 0.100 m mark with a commensurate shift in the other lenses.  The effect of this move was move the first waist (and the position of the EOM back a couple of bolt holes.  The change in size of the first waist was small so the solution remained the same with small tweaks to the second and third lens positions. 

In order to fit the components into the vacuum tank and get a beat note as soon as possible we opted to route the beam along a different route to the north path.  The path is shown with positions of components in the attached schematic. Lenses were mounted on linear translations stages for fine alignment and the Faraday isolator polarization axis was slightly altered to give a return path beam that was parallel to the table.  The way the Thorlabs FI come from the manufacture is such that the input polarization is optimized for horizontal polarization input, rather than reflection in the plane of the beam.

With the changes to the end PBS rotation we re measured the through put power and reverse power attenuation. Out of 9.05 mW incident power the forward propagating output was 8.75 mW: this is a 3.3 % loss. The reverse power throughput was 3.062 uW out of 9.01 mW (we rotated through all input polarization and chose the maximum value): this is a 34.7 dB degree of isolation for back reflection.

A quick measurement of the return path efficiency (of back reflected light extracted from the first FI's first PBS port for PDH locking of the cavity) was 8.07 mW from a total of 9.01 mW or 10.4 % loss.  This was a quick measurement, and maybe something we might like to confirm later. 

---

All the components up the periscope have been aligned with their heights and centering checked. I did a quick walk of the lenses to see if I could find a waist of 220 um located at the expected point for the cavity (checked with a flipper mirror on the table). I ran out of range on the second lens (it needs to be moved back more).  However, the solution is pretty close now. 

 

Also, we need to check the beam waist fitting method (see note PSL_Lab/1680), it may be necessary to do a baseline subtraction of the dark CCD detector to be able to use the D4σ method with confidence.  Otherwise double check the the 1/e^2 clipping option in the WinCamD software; you can get the software to display both at once.  We should compare tight waists to beams much further out so see how the two methods scale.

 

First beatnote is close.

 

-awade Sat Jul 23 01:15:45 2016

Attachment 1: North_South_layout_progress_dist_final_v8.pdf  113 kB  | Hide | Hide all
North_South_layout_progress_dist_final_v8.pdf
ELOG V3.1.3-